# Composite likelihood inference for the Poisson log-normal model

### v1

@@ Abstract
Inferring parameters of a latent variable model can be a daunting task when the conditional distribution of the latent variables given the observed ones is intractable. Variational approaches prove to be computationally efficient but, possibly, lack theoretical guarantees on the estimates, while sampling based solutions are quite the opposite. Starting from already available variational approximations, we define a first Monte Carlo EM algorithm to obtain maximum likelihood estimators, focusing on the Poisson log-normal model which provides a generic framework for the analysis of multivariate count data. We then extend this algorithm to the case of a composite likelihood in order to be able to handle higher dimensional count data.
@@ 1 Introduction
1.1 Maximum likelihood and the EM algorithm
Variational approximation of EM.
1.2 Monte-Carlo approximations of EM
Composite likelihood.
@@ 1.4 Contribution
Our contribution is three-fold. First, we introduce a Monte Carlo EM for the inference of the Poisson log-normal model. Importantly, the proposed algorithm gives access to the asymptotic variance of the parameters and, consequently, to confidence intervals and tests of the parameters, as opposed to the available VEM algorithm. Our second contribution is the use of composite likelihood inference, using blocks of species. This allows to consider latent variables living in a space with smaller dimension (actually, the number of species in each block), so to make importance sampling efficient. To the best of our knowledge, this results in the first EM algorithm for composite likelihood inference in the PLN model framework. Third, we show how to define a mixture proposal distribution that ensures finite variance on bounded test functions and controls the efficiency of the IS steps.
@@ 2.1 Maximum likelihood inference using EM
Asymptotic variance of ML estimates.
@@ 2.2 Variational EM
The VEM alternates the update of the variational parameter ${\boldsymbol{\psi}}$, with this of the model parameter ${\boldsymbol{\theta}}$. The reader may refer to Jaakkola (2001) or Wainwright and Jordan (2008) for a general introduction. An important feature of variational inference is that the resulting estimator of ${\boldsymbol{\theta}}$ does not enjoy the same properties as the regular maximum-likelihood estimator; in particular, no general results ensures its asymptotic normality, nor even its consistency. In summary, the computational efficiency of variational approximation comes at the price of the loss of general statistical guarantees.
@@ 2.3 Composite likelihood inference
Algorithm 2 enjoys the same general property as the EM Algorithm 1. The proof follows the same line as this for EM and is given in Appendix A.1.
@@
In the context of the PLN model, the blocks will consists of subsets of $k$ species, that is $\mathcal{C}_{b}\subset\{1,\dots,p\}$ and ${\mathbf{Y}}^{(b)}=\{Y_{ij}\}_{1\leq i\leq n,j\in\mathcal{C}_{b}}$. We shall consider the same blocks of species for ${\mathbf{Z}}$ as for ${\mathbf{Y}}$, that is ${\mathbf{Z}}^{(b)}=\{Z_{ij}\}_{1\leq i\leq n,j\in\mathcal{C}_{b}}$. Obviously, the conditional moments to be evaluated at the CL-E step are the same than these given in (7), replacing ${\mathbf{Y}}$ (resp. ${\mathbf{Z}}$) with ${\mathbf{Y}}^{(b)}$ (resp. ${\mathbf{Z}}^{(b)}$) for each block $\mathcal{C}_{b}$, but with a further simplification on the parameter. The complete likelihood $p_{\boldsymbol{\theta}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})$ of block $\mathcal{C}_{b}$ does not depend on all the elements of ${\boldsymbol{\theta}}$. More specifically, it only involves ${\boldsymbol{\theta}}^{(b)}=({\bf B}^{(b)},{\boldsymbol{\Sigma}}^{(b)})$ where
${\bf B}^{(b)}=[{\boldsymbol{\beta}}_{j}]_{j\in\mathcal{C}_{b}},\quad\text{and}\quad{\boldsymbol{\Sigma}}^{(b)}=[{\boldsymbol{\Sigma}}_{jk}]_{(j,k)\in\mathcal{C}_{b}}.$
@@
Conversely, the data block ${\mathbf{Y}}^{(b)}$ contributes only to the estimation of the elements ${\boldsymbol{\theta}}^{(b)}$. As a consequence, to be able to estimate each element of the covariance matrix ${\boldsymbol{\Sigma}}$, we need to resort to overlapping blocks, so that each pair of species $(j,k)$ appears at least once in the same block. The construction of the blocks will be discussed in Section 4.1.
Asymptotic variance of MCL estimates
@@
$H({\boldsymbol{\theta}})=\sum_{b=1}^{C}\lambda_{b}\mathbb{E}_{{\boldsymbol{\theta}}}\mathopen{}\mathclose{{}\left[\nabla_{{\boldsymbol{\theta}}}\log p_{{\boldsymbol{\theta}}}({\mathbf{Y}}^{(b)})\{\nabla_{{\boldsymbol{\theta}}}\log p_{{\boldsymbol{\theta}}}({\mathbf{Y}}^{(b)})\}^{\top}}\right].$
@@ 3 Importance Sampling within EM algorithm
3.1 Maximum likelihood inference using IS
@@
3.2 Maximum composite likelihood inference using IS
@@
$\widehat{J}_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\widehat{S}_{i}\widehat{S}_{i}^{\top}-\left\{\frac{1}{n}\sum_{i=1}^{n}\widehat{S}_{i}\right\}\left\{\frac{1}{n}\sum_{i=1}^{n}\widehat{S}_{i}\right\}^{\top}\quad\text{and}\quad\widehat{H}_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\sum_{b=1}^{C}\lambda_{b}\widehat{S}^{(b)}_{i}{\widehat{S}_{i}^{(b)\top}},$ (18)
@@
$\widehat{S}_{i}^{(b)}=\widehat{\mathbb{E}}_{q^{(h)}_{i}}\mathopen{}\mathclose{{}\left[\nabla\log p_{\theta}\left(Y^{(b)}_{i},Z^{(b)}_{i}\right)}\right]\quad\text{and}\quad\widehat{S}_{i}=\sum_{b=1}^{C}\lambda_{b}\widehat{S}_{i}^{(b)}.$
From a numerical perspective, the computation of the former estimators is somewhat less straightforward than those used for the Fisher information matrix. These matrices do not depend on precisely the same statistics as those involved in the M Step, owing to differences in the cross-products utilized. We need the entire weighted sample of the last step to avoid computing unnecessary quantities at each step of the EM procedure.
@@ 3.3 Importance proposal distribution
We choose here to build an initial proposal distribution upon the variational approximation from Chiquet et al. (2018, 2019) and subsequently adapt it at each iteration. Our approach relies on the following result that gives conditions, when the proposal distribution $q_{i}^{(h)}$ is Gaussian, to achieve non-normalized weights $\rho^{(h)}_{i}({\mathbf{V}}_{i})$ that (i) have finite variance and (ii) are bounded. Proposition 3 give conditions under which both properties hold; its proof is given in Appendix A.1.
@@
This choice is surprisingly marred by a major drawback. The computation of the non-normalized weight $\rho^{(h)}_{i}({\mathbf{V}}_{i})$ still involves in the numerator a dominant quadratic term related not to the conditional variance, but to the marginal variance ${\boldsymbol{\Sigma}}^{(h)}$ (see Equation (21) in Appendix A.1). Using a normal proposal based on the conditional variance yields a distribution that is more concentrated than the joint distribution involved in the non-normalized weight. The latter are thus never bounded in the tails. Moreover the condition given by Proposition 3 to get at least finite variance does not hold in all generality. Such a proposal distribution thus provides a fair idea of where to sample, but may nonetheless lead to a zero effective sample size.
@@
where $({\mathbf{m}}^{(0)}_{i},{\mathbf{S}}^{(0)}_{i})$ and ${\boldsymbol{\Sigma}}^{(0)}$ are respectively the variational parameter $({\mathbf{m}}_{i},{\mathbf{S}}_{i})$ and the variational estimation of ${\boldsymbol{\Sigma}}$, and for $h\geq 1$, ${\mathbf{m}}^{(h)}_{i}$ and ${\mathbf{S}}^{(h)}_{i}$ are Monte Carlo estimates of the mean and variance of the conditional distribution $p_{{\boldsymbol{\theta}}^{(h-1)}}(\cdot\;|\;{\mathbf{Y}}_{i})$. The following proposition states that the importance weights associated with the mixture inherit the same properties as the non-normalized weight related to $\varphi(\cdot\,;\,m,{\boldsymbol{\Sigma}}^{(h)})$. The proof follows directly from Proposition 3 and is given in Appendix A.1.
@@
The mixture proportion $\alpha$ controls how close we aim to be from the optimal Gaussian proxy of the conditional, while the second component is solely dedicated to ensuring finite variance of the non-normalized weights in Proposition 4. Algorithm 3 summarizes our MCEM algorithm for the inference of the PLN model (1) based on the mixture proposal 19.
@@ 4 Illustrations
In this section, we first discuss the construction of the blocks in view of composite likelihood inference. Then, we compare and assess the performances of the different inference algorithms previously proposed on synthetic datasets. Eventually, we use them to analyse populations of fishes from the Barents sea.

### v2

@@ Abstract
The Poisson log-normal model is a latent variable model that provides a generic framework for the analysis of multivariate count data. Inferring its parameters can be a daunting task since the conditional distribution of the latent variables given the observed ones is intractable. For this model, variational approaches are the golden standard solution as they prove to be computationally efficient but lack theoretical guarantees on the estimates. Sampling based solutions are quite the opposite. Starting from already available variational approximations, we define a first Monte Carlo EM algorithm to obtain maximum likelihood estimators. We then extend this algorithm to the case of a composite likelihood in order to be able to handle higher dimensional count data.
@@ 1.4 Contribution
Our contribution is three-fold. First, we introduce a Monte Carlo EM for the inference of the Poisson log-normal model. Importantly, the proposed algorithm gives access to the asymptotic variance of the parameters and, consequently, to confidence intervals and tests of the parameters, as opposed to the available VEM algorithm. Our second contribution is the use of composite likelihood inference, using blocks of species. This allows to consider latent variables living in a space with smaller dimension (actually, the number of species in each block), so to make importance sampling efficient. To the best of our knowledge, this results in the first EM algorithm combining importance sampling and composite likelihood to address inference on such a model. Third, we show how to define a mixture proposal distribution that ensures finite variance on bounded test functions and controls the efficiency of the importance sampling steps.
@@ 2.1 Maximum likelihood inference using EM
Asymptotic variance of the maximum likelihood estimates.
@@ 2.2 Variational EM
The VEM alternates the update of the variational parameter ${\boldsymbol{\psi}}$, with this of the model parameter ${\boldsymbol{\theta}}$. The reader may refer to Jaakkola (2001) or Wainwright and Jordan (2008) for a general introduction.
@@
Variance of the variational estimates.
An important feature of variational inference is that the resulting estimator of ${\boldsymbol{\theta}}$ does not enjoy the same properties as the regular maximum-likelihood estimator; in particular, no general results ensures its asymptotic normality, nor even its consistency. In summary, the computational efficiency of variational approximation comes at the price of the loss of general statistical guarantees. In particular, the (asymptotic) variance of the variational estimator is not known. Still, this variance can be estimated using jackknife (Efron and Stein, 1981), which requires to run the VEM algorithm $n$ times. This option is implemented in the PNmodels R package.
@@ 2.3 Composite likelihood inference
Algorithm 2 enjoys the same general property as the EM Algorithm 1. The proof follows the same line as this for EM and is given in Appendix A.1.1.
@@
In the context of the PLN model, the blocks will consists of subsets of $k$ species, that is $\mathcal{C}_{b}\subset\{1,\dots,p\}$ and ${\mathbf{Y}}^{(b)}=\{Y_{ij}\}_{1\leq i\leq n,j\in\mathcal{C}_{b}}$. We shall consider the same blocks of species for ${\mathbf{Z}}$ as for ${\mathbf{Y}}$, that is ${\mathbf{Z}}^{(b)}=\{Z_{ij}\}_{1\leq i\leq n,j\in\mathcal{C}_{b}}$. Obviously, the conditional moments to be evaluated at the CL-E step are the same than these given in (7), replacing ${\mathbf{Y}}$ (resp. ${\mathbf{Z}}$) with ${\mathbf{Y}}^{(b)}$ (resp. ${\mathbf{Z}}^{(b)}$) for each block $\mathcal{C}_{b}$, but with a further simplification on the parameter. The complete likelihood $p_{\boldsymbol{\theta}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})$ of block $\mathcal{C}_{b}$ does not depend on all the elements of ${\boldsymbol{\theta}}$. More specifically, it only involves ${\boldsymbol{\theta}}^{(b)}=({\boldsymbol{B}}^{(b)},{\boldsymbol{\Sigma}}^{(b)})$ where
${\boldsymbol{B}}^{(b)}=[{\boldsymbol{\beta}}_{j}]_{j\in\mathcal{C}_{b}},\quad\text{and}\quad{\boldsymbol{\Sigma}}^{(b)}=[{\boldsymbol{\Sigma}}_{jk}]_{(j,k)\in\mathcal{C}_{b}}.$
@@
Conversely, the data block ${\mathbf{Y}}^{(b)}$ contributes only to the estimation of the elements ${\boldsymbol{\theta}}^{(b)}$. As a consequence, to be able to estimate each element of the covariance matrix ${\boldsymbol{\Sigma}}$, we need to resort to overlapping blocks, so that each pair of species $(j,\ell)$ appears at least once in the same block. The construction of the blocks will be discussed in Section 4.2.
Asymptotic variance of the maximum composite likelihood estimates.
@@
$H({\boldsymbol{\theta}})=\sum_{b=1}^{C}\lambda_{b}\mathbb{E}_{{\boldsymbol{\theta}}}\mathopen{}\mathclose{{}\left[\nabla_{{\boldsymbol{\theta}}}\log p_{{\boldsymbol{\theta}}}({\mathbf{Y}}^{(b)})\{\nabla_{{\boldsymbol{\theta}}}\log p_{{\boldsymbol{\theta}}}({\mathbf{Y}}^{(b)})\}^{\top}}\right]^{\blacksquare}.$
where $A^{\blacksquare}$ stands for a $k\times k$ matrix $A$ that has been embedded into a $p\times p$ matrix $M$ with $M_{j\ell}$ taking value 0 if the pair $(j,\ell)$ does not belong to the possible pair of species of block $\mathcal{C}_{b}$ and being equal to the coefficient of $A$ related to the pair $(j,\ell)$ otherwise.
@@
Model selection.
Model selection procedure also exists in the context of composite likelihood inference and analogues of the AIC (Varin and Vidoni, 2005) and BIC (Gao and Song, 2010) criteria have been derived in this context. The latter reference proves the consistency of the composite likelihood BIC criterion, defined as
$BIC=c\ell_{{\boldsymbol{\theta}}}({\mathbf{Y}})-\frac{\log(n)}{2}\text{dim}({\boldsymbol{\theta}}),\quad\text{where}\quad\text{dim}({\boldsymbol{\theta}})=\mathrm{tr}[H({\boldsymbol{\theta}})G({\boldsymbol{\theta}})^{-1}].$ (14)
Observe that, in the regular likelihood context, both $H({\boldsymbol{\theta}})G({\boldsymbol{\theta}})^{-1}=J({\boldsymbol{\theta}})H({\boldsymbol{\theta}})^{-1}$ reduces to the identity matrix and $\dim({\boldsymbol{\theta}})$ is the number of independent parameters, which yields in the regular BIC formulagit status. In Section 3.2, we will derive estimates of the matrices $H({\boldsymbol{\theta}})$, $J({\boldsymbol{\theta}})$ and $G({\boldsymbol{\theta}})$, which will enable us to evaluate the composite likelihood BIC criterion.
@@ 3 Importance Sampling within EM algorithm
3.1 Maximum likelihood inference using importance sampling
@@
3.2 Maximum composite likelihood inference using importance sampling
@@
$\widehat{J}_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\widehat{S}_{i}\widehat{S}_{i}^{\top}\quad\text{and}\quad\widehat{H}_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\sum_{b=1}^{C}\lambda_{b}\widehat{S}^{(b)}_{i}{\widehat{S}_{i}^{(b)\top}},$ (19)
@@
$\widehat{S}_{i}^{(b)}=\widehat{\mathbb{E}}_{q^{(h)}_{i}}\mathopen{}\mathclose{{}\left[\nabla\log p_{\theta}\left(Y^{(b)}_{i},Z^{(b)}_{i}\right)}\right]^{\blacksquare}\quad\text{and}\quad\widehat{S}_{i}=\sum_{b=1}^{C}\lambda_{b}\widehat{S}_{i}^{(b)},$
where $S^{\blacksquare}$ stands for a vector $S$ of length $d\times k+k(k+1)/2$ that has been embedded into a vector $T$ of length $d\times p+p(p+1)/2$ with $T_{j}$ taking value 0 if the parameter ${\boldsymbol{\theta}}_{j}$ is not involved in the block $\mathcal{C}_{b}$ and being equal to the coefficient of $S$ related to the parameter ${\boldsymbol{\theta}}_{j}$ otherwise. From a numerical perspective, the computation of the former estimators is somewhat less straightforward than those used for the Fisher information matrix. These matrices do not depend on precisely the same statistics as those involved in the M Step, owing to differences in the cross-products utilized. We need to store the statistics specific to each block and each site.
@@ 3.3 Importance proposal distribution
We choose here to build an initial proposal distribution upon the variational approximation from Chiquet et al. (2018, 2019) and subsequently adapt it at each iteration. Our approach relies on the following result that gives conditions, when the proposal distribution $q_{i}^{(h)}$ is Gaussian, to achieve non-normalized weights $\rho^{(h)}_{i}({\mathbf{V}}_{i})$ that (i) have finite variance and (ii) are bounded. Proposition 3 give conditions under which both properties hold; its proof is given in Appendix A.1.2.
@@
This choice is surprisingly marred by a major drawback. The computation of the non-normalized weight $\rho^{(h)}_{i}({\mathbf{V}}_{i})$ still involves in the numerator a dominant quadratic term related not to the conditional variance, but to the marginal variance ${\boldsymbol{\Sigma}}^{(h)}$ (see Equation (23) in Appendix A.1.2). Using a normal proposal based on the conditional variance yields a distribution that is more concentrated than the joint distribution involved in the non-normalized weight. The latter are thus never bounded in the tails. Moreover the condition given by Proposition 3 to get at least finite variance does not hold in all generality. Such a proposal distribution thus provides a fair idea of where to sample, but may nonetheless lead to a zero effective sample size.
@@
where $({\mathbf{m}}^{(0)}_{i},{\mathbf{S}}^{(0)}_{i})$ and ${\boldsymbol{\Sigma}}^{(0)}$ are respectively the variational parameter $({\mathbf{m}}_{i},{\mathbf{S}}_{i})$ and the variational estimation of ${\boldsymbol{\Sigma}}$, and for $h\geq 1$, ${\mathbf{m}}^{(h)}_{i}$ and ${\mathbf{S}}^{(h)}_{i}$ are Monte Carlo estimates of the mean and variance of the conditional distribution $p_{{\boldsymbol{\theta}}^{(h-1)}}(\cdot\;|\;{\mathbf{Y}}_{i})$. The following proposition states that the importance weights associated with the mixture inherit the same properties as the non-normalized weight related to $\varphi(\cdot\,;\,m,{\boldsymbol{\Sigma}}^{(h)})$. The proof follows directly from Proposition 3 and is given in Appendix A.1.2.
@@
The mixture proportion $\alpha$ controls how close we aim to be from the optimal Gaussian proxy of the conditional, while the second component is solely dedicated to ensuring finite variance of the non-normalized weights in Proposition 4. Algorithm 3 summarizes our MCEM algorithm for the inference of the PLN model (1) based on the mixture proposal 20.
@@
Relation to adaptive importance sampling.
Adaptive version of Gaussian mixture proposal distributions can be achieved with the M-PMC algorithm of Cappé et al. (2008). Specifically for the PLN model, in order to update the proposal distribution at each iteration $h$ for each site $i$, and potentially each block $b$, we may consider their defensive option, incorporating a Gaussian distribution with a covariance matrix that fulfils condition (i) of Proposition 3 as defensive component, e.g., for $10\leq\alpha<$ 1 and $K\in\mathbb{N}^{*}$,
$\widetilde{q}^{(h)}_{i}=(1-\alpha)\varphi\left(\cdot\,;\,{\mathbf{m}}^{(h)}_{i},{\boldsymbol{\Sigma}}^{(h)}\right)+\alpha\sum_{k=1}^{K}\eta_{k}\varphi\left(\cdot\,;\,\widetilde{{\mathbf{m}}}^{(h)}_{ik},\widetilde{{\mathbf{S}}}^{(h)}_{ik}\right).$ (21)
If we set the proposal distribution to lie within the family of two-component Gaussian mixtures ($K=1$), the method resume to recursively update the mean and the covariance matrix of the free component to minimizing the Kullback-Leibler divergence between the conditional distribution of interest and the mixture distribution. The latter optimum differs from our solution that solely aims at minimizing the Kullback-Leibler divergence between the conditional distribution of interest and the free component. A numerical comparison of the performances will be presented in Section 4.1.
@@ 4 Illustrations
In this section, we first provide a comparison of our importance sampling setting with an adaptive importance sampling scheme in the context of Gaussian mixtures. Secondly, we discuss the construction of the blocks in view of composite likelihood inference. Then, we compare and assess the performances of the different inference algorithms previously proposed on synthetic datasets. We also present a comparison with the original VEM algorithm. Eventually, we use them to analyse populations of fishes from the Barents sea.
4.1 Comparison with adaptative importance sampling in a Gaussian mixture family
Denote $\widetilde{w}^{(h)}_{ir}$, $1\leq r\leq N$, the importance weights associated with a $N$-sample from the mixture proposal (21) and $\mathcal{R}$ the relative efficiency in terms of the normalized perplexity of the importance weights, that is
$\mathcal{R}=\frac{\exp\left(-\sum_{r=1}^{N}w^{(h)}_{ir}\log w^{(h)}_{ir}\right)}{\exp\left(-\sum_{r=1}^{N}\widetilde{w}^{(h)}_{ir}\log\widetilde{w}^{(h)}_{ir}\right)}.$
Over a data set of 1000 count matrices randomly draws from the PLN model (1) with $n=100$, $p=2,3,5$ and $7$, and $d=3$, Table 1 shows that the adaptive scheme converges towards a mixture for which the Shannon entropy of the associated normalized weight is equivalent to our proposal.
Table 1: Normalized perplexity relative efficiency $\mathcal{R}$ for $N=5000$ draws. Quantiles of the relative efficiency observed over a data set of a 1000 count matrices corresponding to 100 independent draws from the PLN model (1) ($n=100$ sites, $d=3$ covariates) for 10 different random parameter configurations.
Number of species $p$ 2 3 5 7
$5\%$-quantile 0.997 0.994 0.990 0.981
$95\%$-quantile 1.123 1.119 1.067 1.036
For the same data set, whether the number of free mixture component in the adaptive scheme is $K=1$ or $K=2$, we can also observe that the $5\%$ and $95\%$ quantiles of Monte Carlo estimates of the Kullback-Leibler divergence $\mathrm{KL}[\widetilde{q}^{(h)}_{i}\|q^{(h)}_{i}]$ is close to zero (Table 2). In the context of the PLN model, the adaptive scheme applied to Gaussian mixture models proves to be of little use. It implies an additional computational burden and does not lead to significant improvement as the proposal $q^{(h)}_{i}$ is already close to the optimal adaptive version.
Table 2: Kullback-Leibler divergence $\mathrm{KL}[\widetilde{q}^{(h)}_{i}\|q^{(h)}_{i}]$ for $N=5000$ draws and mixture proposal $\widetilde{q}^{(h)}_{i}$ with $K=1$ or $K=2$ free components. Quantiles of the Monte Carlo estimates for a data set of a 1000 count matrices corresponding to 100 independent draws from the PLN model (1) ($n=100$ sites, $d=3$ covariates) for 10 different random parameter configurations.
$K=1$ $K=2$
Number of species $p$ 2 3 5 7 2 3 5 7
$5\%$-quantile (×$\times 10^{-3}$) 4.28 1.77 2.90 5.87 6.27 4.77 7.98 13.7
$95\%$-quantile (×$\times 10^{-1}$) 1.31 1.36 0.80 0.59 1.35 1.38 0.83 0.71